OpenStudy (anonymous):

Solve the system. y = 2 - x and y = 4x + 7 A. (1, 1) B. (1, 3) C. (-1, 3) D. (-1, 1) E. no solution

OpenStudy (anonymous):

C(-1,3)

OpenStudy (anonymous):

x=-1,y=3 so C(-1,3 ) is correct

OpenStudy (anonymous):

Eq. One y=2-x x+y=2 Eq. Two y=4x+7 4x-y=-7 Now Solving two eq.  Now put value of X in eq one y=2-x y=2-(-1) y=2+1 Y=3 So the answer is (-1,3)